What Is The Multiplicative Inverse Of Negative 7
Answer by stanbon(48575)

What Is The Multiplicative Inverse Of Negative 7
--------------
The multiplicative Inverse of ANY number except zero, is 1 over the number.
------------
Your Problem:
MI = 1/(-7) = -1/7
=======================
Cheers,
Stan H.